OpenStudy (anonymous):

NEED HELP my brother is half as old as i am. my sister is twice as old as i am. the sum of our ages is 35. how old am i?

OpenStudy (ajprincess):

Say ur age is 2x. Then ur brother's age will be x. Also ur sister's age will be 4x. 2x+x+4x=35 solve x. Then find 2x nd that will be your age.

OpenStudy (ajprincess):

Can u do it @arlinda27?

OpenStudy (lgbasallote):

immortal

OpenStudy (anonymous):

Let, x = your age y = brother's age z = sister's age. As your brother age is half of yours, i.e. y = x/2 Your sister is twice as old as you, i.e. z = 2x = 4y Now, the sum of your age = 35 or, x + y + z = 35 or, 2y + y + 4y = 35 or y = 5 and x = 2 * 5 = 10 Ans.

OpenStudy (anonymous):

thank u all

OpenStudy (anonymous):

you are most welcomed.